In its first year, “Helen Trading” had the following experience:

















Unit selling price birr 100 Sales 50,000 units

















Total fixed cost birr 400,000 TVC birr 3,000,000

















Required:

















a) Develop TR, TC and profit functions for the company in terms of quantity

















b) Find the breakeven point in terms of quantity and sales volume.

















c) If profit had been br. 500,000 what would have been the sales volume (Revenue) and quantity of sales.

















d) What would have been the profit if sales are br. 2,000,000

















e) Draw Breakeven chart and discuss about it














1
Expert's answer
2022-02-24T11:38:36-0500

a)

Total Revenue function

= P"\times" Q

P is given as 100

= 100Q is revenue function


Total cost function

= FC + VC

= 400000 + ( "\frac{3000000} {50000})" "\times" Q

= 400000 + 60Q


Profit function

= TR - TC

= 100Q - 400000 - 60Q

= 40Q - 400000



b)Find the break even point in terms of quantity and sales volume


Break even quantity

= "\frac{Fixed cost }{( P - VC )}"

= "\frac{400000} {( 100 - 60 )}"

= 10000 units is break even quantity


Sales volume

= P "\times" Q

= 100 "\times" 10000

= 1000000 is sales volume in terms of revenue



c)If profit had been 500000 what would have been the sales volume (Revenue) and quantity of sales.


Profit

Profit = 40Q - 400000

500000 = 40Q - 400000

40Q = 900000

Q = 22500 units require to get profit of 500000


Sales revenue

= P "\times" Q

= 100 "\times" 22500

= 2250000


d)What would have been the profit if sales are 2000000

Sales revenue = 2000000

Quantity = "\frac{Revenue} {Price}"

Quantity = "\frac{2000000} { 100}" = 20000 units


Profit = 40Q - 400000

Profit = 40"\times" 20000 - 400000

Profit = 800000 - 400000

Profit = 400000 is profit

e)